Faraday’s Law

1. Hydrogen fuel cells are based on the reaction of hydrogen and oxygen to form water:

2H 2 (g) + O 2 (g) ? 2H 2O (l)

In the presence of excess oxygen, a hydrogen fuel cell consumes 0.370 grams of hydrogen per hour at 25°C. Calculate the maximum current that can be produced by

this fuel cell.

2. Aluminum is produced by the electrolytic reduction of alumina, Al 2O 3. The overall cell reaction can be written as:

2Al 2O 3 (s) + 3C (s) ? 4Al (s) + 3CO 2 (g)

a) Given a current of 1.0 Å~ 10 5 amps, how many hours would it take to produce 1000 kg of aluminum?

b) The anode in the above reaction is graphite (carbon), which is oxidized to CO 2 (g) during the reaction. What mass of graphite must be consumed in order to produce 1000 kg of aluminum?
Answer #1

1)

2H2(g) + O2 ----> 2H2O

according to Faraday first law W = z* C * t

w = mass of the substance deposited

Z = electro chemical equivalent of substance

C = current passed through the circuit

t = time taken for the current passed = 1 hour

0.37 = 2/96500*C*60*60

C = 4.96 amp

2)

2Al2O3 + 3c ----> 4 Al + 3CO2

according to Faraday first law,

w = z * C * t

1000*10^3 = (27/3)/96500 * 10^5 * t

t = 107222.22 sec

t = 107222.22/(60*60)

t = 29.783 hrs

b)

using first law of Faraday,

W = z * C * t

w = (12/4)/96500*10^5*107222.222

w = 333.333 kg of graphite is used.